OpenStudy (anonymous):

A system of equations is shown below. 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Part A) "Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other." We will replace the first equation by multiplying the second equation by, say 2, and adding it to the first equation like the instruction asks us to do: 2x + 2y = 3 multiply by 2 4x + 4y = 6 add it to the first equation: 4x + 4y = 6 3x + 8y = 12 add: 7x + 12y = 18 So the new system of equations are: 7x + 12y = 18 2x + 2y = 3 For part B) solve the original system of equations for x and y. Then solve the system of equations created in part A for x and y And show that you get the same answers.

OpenStudy (anonymous):

@ranga solving for x and y and soling the originl system of equations is what i need help with